A certain radioactive element A has a half life =t seconds.In t/2 seconds the fraction of the initial quantity of the element so far decayed is nearly

Grade:12

2 Answers

Aarti Gupta
10 years ago
Half - life= t seconds,it means if X gm is taken initially then X/2 gm remains in t seconds.Thus in t/2 seconds,fraction decayed will be X/4 gm,or 25% will decayed.
Thanks & Regards
Aarti Gupta
askiitians Faculty
Dharmendra soni
4 years ago
K=0.693/t
Kt=ln(A1/A2)
(0.693/t)*t/2=ln(A1/A2)
.693/2=ln(A1/A2)
A1/A2=e^(.3465).    .....    (e^. 3465=1.414)
A2/A1=1/1.414
A2/A1=0.7072
(in %) 
70.72%(is remained) 
It means it decayed 29.28%(answer)